OpenStudy (anonymous):

Answer this: (1+2(3+4))/(5*6)

OpenStudy (anonymous):

(1+2(7))/30 =(1+14)/30 =15/30 =1/2

OpenStudy (anonymous):

why did you multiply 2 and 7

OpenStudy (anonymous):

3+4 is 7. 2(7) is the same as 2*7

OpenStudy (anonymous):

2(3+4) =2(7) =14

OpenStudy (anonymous):

ohh okay thanks

OpenStudy (anonymous):

yeah no problem